OpenStudy (anonymous):

The sides of an isosceles triangle have lengths 2x + 4, x + 8. The base has length 5x − 2. What is the length of the base?

hartnn (hartnn):

so you know that the 2 sides in an isosceles triangle are equal? so we can do, \(2x+4 = x+8\) Try to find x from this?

OpenStudy (anonymous):

x= 4

hartnn (hartnn):

thats correct! :) now just plug x = 4 in 5x-2 \(5(4) - 2 = ..?\) that would be the length ofbase.

OpenStudy (anonymous):

thanks ;)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Two sides of an isosceles triangle are equal. Using the above statement I may form the following equation: 2x + 4 = x + 8 Now subtract x on both sides i.e, \[2x - x + 4 = x + 8 - x\] Now, 2x - x comes down to just x so the LHS is : x+4 For the RHS , - x cancels out + x Therefore, RHS = 8 Now equate the new LHS to the new RHS that is: x + 4 = 8 This time subtract 4 on both sides. Again, by repeat the above steps i.e, \[x + 4 - 4 = 8- 4\]

hartnn (hartnn):

nice explanation @MatthewPFDS :)

OpenStudy (anonymous):

It seems someone else already explained it to you.. and thanks @hartnn :)